# A model for Axial Non-Standard Interactions of neutrinos with quarks

### v1

@@ IV Possible generalization of the toy model
In sect. II, we presented a model for $\epsilon_{\tau\tau}^{Au}=\epsilon^{Ad}_{\tau\tau}$. The reason why we have focused on the $\tau\tau$ component is that this component is less experimentally restricted. We could also assign a $U(1)$ charge to the second generation of leptons and obtain $\epsilon_{\mu\mu}^{Au}=\epsilon^{Ad}_{\mu\mu}\neq 0$. As long as the $U(1)$ charges of the second and third generation leptons summed up to 9, the $U(1)-SU(2)-SU(2)$ anomaly would be canceled. Exploring the parameter space is beyond the scope of this paper. Similarly to the models for vectorial NSI presented in Farzan (2020); Farzan and Heeck (2016), arbitrary flavor structure can be obtained by adding sterile neutrinos with a $U(1)$ charge mixed with the active neutrinos.
To cancel the $U(1)-SU(3)-SU(3)$ anomaly, we have added a pair of quarks. The lower bound on the masses of these colored particles leads to an upper bound on $\epsilon^{Au}=\epsilon^{Ad}$. Instead of introducing new quarks, we could assign $U(1)$ charges to $s_{R}$ and $c_{R}$, $\alpha_{s}$ and $\alpha_{c}$, such that $\alpha_{s}+\alpha_{c}=6$. This requires new Higgs doublets with suitable $U(1)$ charges to give mass to the first and second quarks. Then, to cancel various $U(1)$ and $U_{Y}(1)$ anomalies, new charged particles (charged both under $U_{em}(1)$ and new $U(1)$) should be added. In the minimal version, with addition of only a pair of chiral fermions, $E_{L}$ and $E_{R}$, with electric charge of 1 and different $U(1)$ charges, the anomaly cancellation requires fractional $U(1)$ charges. The mass of $E$ should be obtained after breaking of the $U(1)$ symmetry via a Yukawa coupling of form $\phi E_{L}^{\dagger}E_{R}$. The new fermion is colorless so the the lower bounds on its mass is less stringent. As a result, $\langle\phi\rangle$ can be smaller than the case with colorful new fermions. Thus, $\epsilon^{Au}=\epsilon^{Ad}$ can be larger in this version of the model without new particles.
@@ V Summary and discussion
To obtain axial coupling, the $U(1)$ charges of right-handed and left-handed first generation quarks have to be opposite which leads to a $U(1)-SU(3)-SU(3)$ anomaly. To cancel the anomaly, we introduce new quarks with masses of $\sim 1500$ GeV. In our model, the mass generation mechanism for the light quarks is not Yukawa coupling to the SM Higgs but mixings with new heavy quarks. In other words, there will be a seesaw mechanism, explaining the smallness of the $u$ and $d$ quark masses relative to the top mass. The heavy quarks can be pair produced at the hadron colliders and can then decay to the first generation quarks plus $W$ or $Z$ gauge boson. Being only slightly heavier than the present lower bound, there is a bright prospect of discovery of the heavy quarks at the near future by the LHC.
Finally, to cancel the $U(1)$ anomaly, a SM singlet chiral fermion, $\chi_{L}$, has to be added. Imposing a $Z_{2}$ symmetry, $\chi_{L}$ can be made stable and therefore a suitable dark matter particle. The dark matter can undergo spin dependent scattering off quarks. The bounds from spin dependent direct dark matter searches already push the $\chi_{L}$ mass to be smaller than 2 GeV. The future spin dependent dark matter experiments can discover $\chi_{L}$. We have discussed the $\chi_{L}$ production mechanism in the early universe and pointed out three alternative scenarios to obtain the observed dark matter abundance. In our model, the indirect dark matter searches are expected to yield null result and the dark matter will act as collision-less at structure formation as well as at collision of the galaxy clusters.
@@ Acknowledgements
This work has partially been supported by the European Unions Framework Programme for Research and Innovation Horizon 2020 under grant H2020-MSCA-ITN2019/860881-HIDDeN as well as under the Marie Sklodowska-Curie Staff Exchange grant agreement No 5 101086085-ASYMMETRY. YF would like to acknowledge support from ICTP through the Associates Programme and from the Simons Foundation through grant number 284558FY19. YF thanks ICTP high energy group for their hospitality during her visit when this paper was written.

### v2

@@ IV Possible generalization of the toy model
In sect. II, we presented a model for $\epsilon_{\tau\tau}^{Au}=\epsilon^{Ad}_{\tau\tau}$. The reason why we have focused on the $\tau\tau$ component is that this component is less experimentally restricted. We could also assign a $U(1)$ charge to the second generation of leptons and obtain $\epsilon_{\mu\mu}^{Au}=\epsilon^{Ad}_{\mu\mu}\neq 0$. As long as the $U(1)$ charges of the second and third generation leptons summed up to 9, the U(1)SU(2)SU(2)U(1)-SU(2)-SU(2) anomaly would be canceled. Exploring the parameter space is beyond the scope of this paper. Similarly to the models for vector NSI presented in Farzan (2020); Farzan and Heeck (2016), arbitrary flavor structure can be obtained by adding sterile neutrinos with a $U(1)$ charge mixed with the active neutrinos.
To cancel the U(1)SU(3)SU(3)U(1)-SU(3)-SU(3) anomaly, we have added a pair of quarks. The lower bound on the masses of these colored particles leads to an upper bound on $\epsilon^{Au}=\epsilon^{Ad}$. Instead of introducing new quarks, we could assign $U(1)$ charges to $s_{R}$ and $c_{R}$, $\alpha_{s}$ and $\alpha_{c}$, such that $\alpha_{s}+\alpha_{c}=6$. This requires new Higgs doublets with suitable $U(1)$ charges to give mass to the first and second quarks. Then, to cancel various $U(1)$ and $U_{Y}(1)$ anomalies, new charged particles (charged both under $U_{em}(1)$ and new $U(1)$) should be added. In the minimal version, with the addition of only a pair of chiral fermions, $E_{L}$ and $E_{R}$, with an electric charge of 1 and different $U(1)$ charges, the anomaly cancellation requires fractional $U(1)$ charges.
The mass of $E$ should be obtained after breaking of the $U(1)$ symmetry via a Yukawa coupling of form $\phi\bar{E}_{L}E_{R}$. The new fermion is colorless so the lower bounds on its mass are less stringent. As a result, $\langle\phi\rangle$ can be smaller than the case with colorful new fermions. Thus, $\epsilon^{Au}=\epsilon^{Ad}$ can be larger in this version of the model without new particles.
@@ V Summary and discussion
To obtain axial coupling, the $U(1)$ charges of right-handed and left-handed first generation quarks have to be opposite which leads to a U(1)SU(3)SU(3)U(1)-SU(3)-SU(3) anomaly. To cancel the anomaly, we introduce new quarks with masses of $\sim 1500$ GeV. In our model, the mass generation mechanism for the light quarks is not Yukawa coupling to the SM Higgs but mixings with new heavy quarks. In other words, there will be a seesaw mechanism, explaining the smallness of the $u$ and $d$ quark masses relative to the top mass. The heavy quarks can be pair produced at the hadron colliders and can then decay to the first generation quarks plus a $W$ or $Z$ gauge boson. Being only slightly heavier than the present lower bound, there is a bright prospect of discovery of the heavy quarks at the near future by the LHC.
Finally, to cancel the $U(1)$ anomaly, a SM singlet chiral fermion, $\chi_{L}$, has to be added. Imposing a $Z_{2}$ symmetry, $\chi_{L}$ can be made stable and therefore a suitable dark matter particle. The dark matter can undergo spin dependent scattering off quarks. The bounds from spin dependent direct dark matter searches already push the $\chi_{L}$ mass to be smaller than 2 GeV. The future spin dependent dark matter experiments can discover $\chi_{L}$. We have discussed the $\chi_{L}$ production mechanism in the early universe and pointed out three alternative scenarios to obtain the observed dark matter abundance. In our model, the indirect dark matter searches are expected to yield null results and the dark matter will act as collision-less at structure formation as well as at collision of the galaxy clusters.
@@ Acknowledgements
This work is based on research funded by Iran National Science Foundation (INSF) under project No.4031487. This work has partially been supported by the European Unions Framework Programme for Research and Innovation Horizon 2020 under grant H2020-MSCA-ITN2019/860881-HIDDeN as well as under the Marie Sklodowska-Curie Staff Exchange grant agreement No 5 101086085-ASYMMETRY. YF would like to acknowledge support from ICTP through the Associates Programme and from the Simons Foundation through grant number 284558FY19. YF thanks the ICTP high energy group for their hospitality during her visit when this paper was written.